# FUSAR-KLIP: Towards Multimodal Foundation Models for Remote Sensing

### v3

@@ 4.3 Visual Task Benchmarks
All models use the visual encoder weights in the multimodal foundation model as initialization parameters for subsequent SFT. In addition, three single-modal VIT foundation models are also used for experimental comparison. Among them, MAE[he2022masked] and MOCO v3[Chen_2021_ICCV] are pre-trained based on the dataset proposed in this paper, while SAR-JEPA[LI2024326] is the foundation model proposed by Li et al. for SAR image interpretation. Table III summarizes the key model configurations and training hyperparameters of each subtask. The training and test data are divided in a ratio of 4:1 by default.
@@
Classification ViT-Cls[dosovitskiy2021an] AdamW 0.003 100 Warmup,CosineDecay
Detection ViTDET[li2022exploring] AdamW 0.0001 36 Warmup,MultiStep
Segmentation Segmenter[strudel2021segmenter] AdamW 0.001 100 PolyLR
@@
1) Target Classification: The classifier is implemented using the PyTorch framework and evaluated on three representative SAR datasets: FUSAR-SHIP [hou2020fusar], FUSAR-AIR [Qian2024Thesis], and SAR-ACD [9754573]. FUSAR-SHIP and FUSAR-AIR are curated subsets of FUSAR-GEOVL-1M, focusing on ship and aircraft target recognition, respectively. FUSAR-SHIP, developed by the Key Laboratory of Electromagnetic Wave Information Science at Fudan University, includes 5,242 images spanning 15 ship categories and 98 subcategories. FUSAR-AIR comprises diverse aircraft types such as transport, refueling, and civil aircraft. SAR-ACD, a public third-party dataset with higher scene complexity and 1-meter resolution, contains 3,032 images across six aircraft types, with approximately 500 samples per class.
@@
2) Target Detection: This study developed an object detection model based on the typical ViT detection framework, ViTDet. By combining the ViT backbone with the feature pyramid structure, ViTDet enhances the detection of multi-scale targets while maintaining global modeling capabilities, achieving excellent performance across various visual detection tasks. We evaluated the model’s performance using three representative SAR image datasets, under the AP@50 metric. FUSAR-SHIP-Sense (FU-SS) consists of FUSAR-SHIP target slices expanded into complete scene images, totaling 3,838 images. FUSAR-AIR-Sense (FU-AS) is composed of FUSAR-AIR target slices expanded into complete scene images, totaling 2,491 images. SAR-AIRcraft-Few (AIR-F)[Wang2023SARAIRcraft] is an open-source aircraft detection dataset built from Gaofen-3 images, containing 4,368 images, 16,463 targets, and 7 aircraft categories. This experiment used only 20% of the training data to assess the generalization ability of the pre-trained model in non-homologous and few-sample scenarios.
@@
3) Segmentation: This study constructed a SAR image semantic segmentation model based on PyTorch, employing ViT as the visual encoder and the Mask Transformer from Segmenter as the decoder to enhance semantic modeling. The model was evaluated using overall accuracy (OA) and mean Intersection over Union (mIoU) on two representative datasets. FUSAR-MAP[shi2021object] focuses on fine-grained segmentation of urban scenes—including buildings, roads, vegetation, and background—with 600 high-resolution SAR images (1024×1024). AIR-PolarSAR-Seg (PoSAR-Seg)[wang2022air] targets six land cover categories across 2,000 image slices (512×512), enabling evaluation of general semantic segmentation performance.
@@ 4.4 Visual-Language Tasks Benchmarks
The experimental configuration is based on the fine-tuning settings of mainstream multimodal models, such as BLIP[li2022blip], used in the natural image domain, with adjustments made to suit the characteristics of the data. Table V summarizes the core training parameters for the three visual language tasks. Considering task-specific differences, the captioning task is configured with a larger maximum token length, while the VQA task uses a slightly higher learning rate to accelerate convergence.
@@
TABLE VI: Benchmarks for visual language tasks, including image-text retrieval, image captioning, and six question-answering tasks. Target Pos: Target Position. Main Land: Main Landforms. All Land: All Landforms. Reg Land: Regional Landforms. Suboptimal results are underlined.
@@
1) Image-Text Retrieval: This task evaluates the model’s cross-modal retrieval capability between SAR images and text in both directions: Image→Text and Text→Image, which is crucial for remote sensing image retrieval. Standard $Recall@K$ metrics ($R@1$, $R@5$, $R@10$) are used to assess performance, with $txt\_{r1/r5/r10}$ and img_r1/r5/r10img\_r1/r5/r10 denoting the Top-K recall rates for text-to-image and image-to-text retrieval, respectively. The average of these six values serves as the overall retrieval metric. Quantitative and qualitative results are shown in Table VI and Fig. 14. As shown in Fig. 14, we compare the proposed method with a suboptimal model. FUSAR-KLIP has the highest matching degree for the correct option and can therefore retrieve the corresponding image more accurately.
2) Image Captioning: The Caption task aims to generate accurate descriptions of key targets and their spatial distribution in SAR images using natural language. We evaluate the quality of the image descriptions using four mainstream metrics[li2022blip]: BLEU-4, METEOR, CIDEr, and SPICE. These metrics assess the description quality from four dimensions: phrase matching, semantic consistency, content relevance, and structural integrity. Table VI presents the quantitative evaluation results for each pre-trained model on the Caption task. FUSAR-KLIP achieves leading performance across many indicators, particularly demonstrating significant advantages in CIDEr.
Table VI summarizes the performance across all six tasks, while Fig. 15 presents qualitative examples of model predictions for representative questions. Experimental results show that FUSAR-KLIP consistently achieves strong performance, especially in target location reasoning and comprehensive landform identification, highlighting its robust reasoning ability in SAR-based visual question answering scenarios.
4.5 Comparison with general multimodal large models
Some commercial multimodal large models with ultra-large parameter volumes (such as GPT-5, Gemini 3, and Grok-4) have demonstrated strong multi-task capabilities. However, due to their closed-source nature, their applicability in the remote sensing field still requires further verification. Given that VQA is the most common application form for multimodal language models, we use the VQA task to assess the reasoning ability of these models on SAR images. The experimental results are shown in Table VII. The general large models perform the weakest in the target counting task, and their accuracy in other tasks is significantly lower compared to the model proposed in this study. This indicates that general multimodal large models suffer from insufficient adaptability in SAR tasks, while the foundational model developed in this research provides robust support for advancing multimodal models in the SAR domain.
TABLE VII: Comparison with General multimodal LLM. Target Rec: Target Recognition. Target Pos: Target Position. Main Land: Main Landforms. All Land: All Landforms. Reg Land: Regional Landforms.
LLMs Target Count Target Rec Target Pos Main Land ALL Land Reg Land
GPT-5[OpenAIChatGPT] 35 67 28 52 49 47
Gemini-3[google_gemini_web] 21 40 41 56 71 42
Grok-4[xai_grok4] 19 29 54 36 67 48
Qwen3-VL[bai2025qwen3vltechnicalreport] 38 51 27 54 34 40
FUSAR-KLIP 98 96 97 99 89 93
@@ 4.6 Ablation experiment
In the data validation, we selected the latest large-scale SAR text dataset, SARLANG[wei2025sarlang], for comparison. This dataset depends on target detection annotations and fixed templates to generate image-text pairs, ensuring high semantic accuracy without the need for additional screening or reconstruction. However, as analyzed in Fig. 11, SARLANG exhibits issues such as sparse semantics and limited knowledge coverage.
@@
To validate the effectiveness of our proposed semantic enhancement strategy, we design a structured evaluation around the screen–filter–reconstruct mechanism, which aims to eliminate low-quality samples and refine textual content to enhance semantic quality and training robustness. Specifically, we conduct two ablation experiments: (1) removing the entire screen–filter–reconstruct pipeline and training directly on the original data; and (2) applying only the screen–filter stage to retain high-quality samples without performing textual reconstruction.
@@ 5 Conclusion
This paper addresses the fundamental cognitive disconnect between general visual representations and remote sensing SAR interpretation logic in cross-modal artificial intelligence, proposing FUSAR-KLIP—the first knowledge-guided general multimodal SAR foundational model. To this end, we constructed FUSAR-GEOVL-1M, the largest and first large-scale SAR image dataset to date that fully integrates geographic projection attributes. Compared to traditional datasets, FUSAR-GEOVL-1M exhibits significant advantages in spatial scale consistency and cross-platform adaptability. Furthermore, through our proposed Hierarchical Cognitive Chain and Self-Consistent Iterative Optimization mechanism, it successfully transforms implicit geographic environment and scattering features into explicit structured knowledge, providing systematic physical and geographic semantic support for cross-modal alignment.
We established a unified evaluation benchmark across 11 typical downstream tasks and conducted a systematic comparison with 15 mainstream multimodal models. Experimental results show that FUSAR-KLIP demonstrates leading performance across multiple tasks, particularly achieving breakthrough improvements in target counting and land cover classification tasks that heavily rely on spatial reasoning. This result strongly validates that incorporating geographical priors and physical cognition into representation learning is a key path to overcoming the bottlenecks in remote sensing interpretation and bridging the gap between machine representation and human cognition.
This study establishes for the first time the core role of ”knowledge-guided” in the construction of remote sensing foundation models and proposes a new paradigm that integrates geographic priors and cognitive modeling. It also provides theoretical reference for modeling other remote sensing modalities with significant physical characteristics, such as multispectral and infrared. In the future, we will focus on expanding data scale and task types, further exploring unified modeling across modalities and domains, and moving towards building a more complete and logically consistent intelligent system for Earth observation.

### v4

@@ 4.3 Visual Task Benchmarks
All models use the visual encoder weights in the multimodal foundation model as initialization parameters for subsequent SFT. In addition, three single-modal VIT foundation models are also used for experimental comparison. Among them, MAE[14] and MOCO v3[7] are pre-trained based on the dataset proposed in this paper, while SAR-JEPA[30] is the foundation model proposed by Li et al. for SAR image interpretation. Table III summarizes the key model configurations and training hyperparameters of each subtask. The training and test data are divided in a ratio of 4:1 by default.
@@
Classification ViT-Cls[12] AdamW 0.003 100 Warmup,CosineDecay
Detection ViTDET[31] AdamW 0.0001 36 Warmup,MultiStep
Segmentation Segmenter[55] AdamW 0.001 100 PolyLR
@@
1) Target Classification: The classifier is implemented using the PyTorch framework and evaluated on three representative SAR datasets: FUSAR-SHIP [16], FUSAR-AIR [49], and SAR-ACD [57]. FUSAR-SHIP and FUSAR-AIR are curated subsets of FUSAR-GEOVL-1M, focusing on ship and aircraft target recognition, respectively. FUSAR-SHIP, developed by the Key Laboratory of Electromagnetic Wave Information Science at Fudan University, includes 5,242 images spanning 15 ship categories and 98 subcategories. FUSAR-AIR comprises diverse aircraft types such as transport, refueling, and civil aircraft. SAR-ACD, a public third-party dataset with higher scene complexity and 1-meter resolution, contains 3,032 images across six aircraft types, with approximately 500 samples per class.
@@
2) Target Detection: This study developed an object detection model based on the typical ViT detection framework, ViTDet. By combining the ViT backbone with the feature pyramid structure, ViTDet enhances the detection of multi-scale targets while maintaining global modeling capabilities, achieving excellent performance across various visual detection tasks. We evaluated the model’s performance using three representative SAR image datasets, under the AP@50 metric. FUSAR-SHIP-Sense (FU-SS) consists of FUSAR-SHIP target slices expanded into complete scene images, totaling 3,838 images. FUSAR-AIR-Sense (FU-AS) is composed of FUSAR-AIR target slices expanded into complete scene images, totaling 2,491 images. SAR-AIRcraft-Few (AIR-F)[61] is an open-source aircraft detection dataset built from Gaofen-3 images, containing 4,368 images, 16,463 targets, and 7 aircraft categories. This experiment used only 20% of the training data to assess the generalization ability of the pre-trained model in non-homologous and few-sample scenarios.
@@
3) Segmentation: This study constructed a SAR image semantic segmentation model based on PyTorch, employing ViT as the visual encoder and the Mask Transformer from Segmenter as the decoder to enhance semantic modeling. The model was evaluated using overall accuracy (OA) and mean Intersection over Union (mIoU) on two representative datasets. FUSAR-MAP[52] focuses on fine-grained segmentation of urban scenes—including buildings, roads, vegetation, and background—with 600 high-resolution SAR images (1024×1024). AIR-PolarSAR-Seg (PoSAR-Seg)[63] targets six land cover categories across 2,000 image slices (512×512), enabling evaluation of general semantic segmentation performance.
@@ 4.4 Visual-Language Tasks Benchmarks
The experimental configuration is based on the fine-tuning settings of mainstream multimodal models, such as BLIP[26], used in the natural image domain, with adjustments made to suit the characteristics of the data. Table V summarizes the core training parameters for the three visual language tasks. Considering task-specific differences, the captioning task is configured with a larger maximum token length, while the VQA task uses a slightly higher learning rate to accelerate convergence.
@@
1) Image-Text Retrieval: This task evaluates the model’s cross-modal retrieval capability between SAR images and text in both directions: Image→Text and Text→Image, which is crucial for remote sensing image retrieval. Standard $Recall@K$ metrics ($R@1$, $R@5$, $R@10$) are used to assess performance, with $txt\_{r1/r5/r10}$ and img_r1/r5/r10img\_r1/r5/r10 denoting the Top-K recall rates for text-to-image and image-to-text retrieval, respectively. The average of these six values serves as the overall retrieval metric. Quantitative and qualitative results are shown in Table VII and Fig. 14. As shown in Fig. 14, we compare the proposed method with a suboptimal model. FUSAR-KLIP has the highest matching degree for the correct option and can therefore retrieve the corresponding image more accurately.
2) Image Captioning: The Caption task aims to generate accurate descriptions of key targets and their spatial distribution in SAR images using natural language. We evaluate the quality of the image descriptions using four mainstream metrics[26]: BLEU-4, METEOR, CIDEr, and SPICE. These metrics assess the description quality from four dimensions: phrase matching, semantic consistency, content relevance, and structural integrity. Table VII presents the quantitative evaluation results for each pre-trained model on the Caption task. FUSAR-KLIP achieves leading performance across many indicators, particularly demonstrating significant advantages in CIDEr.
@@
Table VII summarizes the performance across all six tasks, while Fig. 15 presents qualitative examples of model predictions for representative questions. Experimental results show that FUSAR-KLIP consistently achieves strong performance, especially in target location reasoning and comprehensive landform identification, highlighting its robust reasoning ability in SAR-based visual question answering scenarios.
4.5 Comparison with general multimodal large models
TABLE VI: Comparison with General multimodal LLM. Target Rec: Target Recognition. Target Pos: Target Position. Main Land: Main Landforms. All Land: All Landforms. Reg Land: Regional Landforms.
LLMs Target Count Target Rec Target Pos Main Land ALL Land Reg Land
GPT-5[46] 35 67 28 52 49 47
Gemini-3[13] 21 40 41 56 71 42
Grok-4[66] 19 29 54 36 67 48
Qwen3-VL[2] 38 51 27 54 34 40
FUSAR-KLIP 98 96 97 99 89 93
TABLE VII: Benchmarks for visual language tasks, including image-text retrieval, image captioning, and six question-answering tasks. Target Pos: Target Position. Main Land: Main Landforms. All Land: All Landforms. Reg Land: Regional Landforms. Suboptimal results are underlined.
@@ 4.5 Comparison with general multimodal large models
Some commercial multimodal large models with ultra-large parameter volumes (such as GPT-5, Gemini 3, and Grok-4) have demonstrated strong multi-task capabilities. However, due to their closed-source nature, their applicability in the remote sensing field still requires further verification. Given that VQA is the most common application form for multimodal language models, we use the VQA task to assess the reasoning ability of these models on SAR images. The experimental results are shown in Table VI. The general large models perform the weakest in the target counting task, and their accuracy in other tasks is significantly lower compared to the model proposed in this study. This indicates that general multimodal large models suffer from insufficient adaptability in SAR tasks, while the foundational model developed in this research provides robust support for advancing multimodal models in the SAR domain.
@@ 4.6 Ablation experiment
In the data validation, we selected the latest large-scale SAR text dataset, SARLANG[64], for comparison. This dataset depends on target detection annotations and fixed templates to generate image-text pairs, ensuring high semantic accuracy without the need for additional screening or reconstruction. However, as analyzed in Fig. 11, SARLANG exhibits issues such as sparse semantics and limited knowledge coverage.
To validate the effectiveness of our proposed SCIO strategy, we design a structured evaluation around the screen–filter–reconstruct mechanism, which aims to eliminate low-quality samples and refine textual content to enhance semantic quality and training robustness. Specifically, we conduct two ablation experiments: (1) removing the entire screen–filter–reconstruct pipeline and training directly on the original data; and (2) applying only the screen–filter stage to retain high-quality samples without performing textual reconstruction.
@@ 5 Conclusion
This paper proposes FUSAR-KLIP, the first knowledge-guided multimodal foundational model for SAR, addressing the cognitive gap between general vision and geoscientific interpretation. Benchmarks across 11 downstream tasks demonstrate that FUSAR-KLIP consistently outperforms 15 mainstream models, particularly in reasoning-heavy tasks like target counting. This result strongly validates that incorporating geographical priors and physical cognition into representation learning is a key path to overcoming the bottlenecks in remote sensing interpretation and bridging the gap between machine representation and human cognition.
More broadly, this work highlights the importance of knowledge-guided modeling in the development of remote sensing foundation models and suggests a unified perspective that integrates geographic priors with cognitive modeling. The proposed framework also offers methodological insights for other remote sensing modalities with strong physical characteristics, such as multispectral and infrared imagery. Future work will focus on expanding data scale and task diversity, as well as further exploring unified modeling across modalities and domains.